OpenStudy (anonymous):

WILL FAN AND MEDAL! :) If f(x) = 2x2 + 4, find f(3). 22 16 -0.5 40

TheSmartOne (thesmartone):

Just plug in 3 in for x.

TheSmartOne (thesmartone):

So \(f(3)=2(3)^2+4\)

TheSmartOne (thesmartone):

Do you know what \(3^2=?\)

TheSmartOne (thesmartone):

@domisreallytall

OpenStudy (anonymous):

9

TheSmartOne (thesmartone):

Yes and \(9\times2=?\)

OpenStudy (anonymous):

18

OpenStudy (anonymous):

22

OpenStudy (anonymous):

f(3)=22

TheSmartOne (thesmartone):

And finally \[18+4=?\]

OpenStudy (anonymous):

thanks !

TheSmartOne (thesmartone):

So yes. f(3)=22 Good Job and thanks for participating.

OpenStudy (anonymous):

but dont you need to find the value of your variable, f,?

TheSmartOne (thesmartone):

no.. f(x) is nothing but y